# Video-Panda: Parameter-efficient Alignment for Encoder-free Video-Language Models

### v1

@@ Abstract
We present an efficient encoder-free approach for video-language understanding that achieves competitive performance while significantly reducing computational overhead. Current video-language models typically rely on heavyweight image encoders (300M-1.1B parameters) or video encoders (1B-1.4B parameters), creating a substantial computational burden when processing multi-frame videos. Our method introduces a novel Spatio-Temporal Alignment Block (STAB) that directly processes video inputs without requiring pre-trained encoders while using only 45M parameters for visual processing - at least a 6.5$\times$ reduction compared to traditional approaches. The STAB architecture combines Local Spatio-Temporal Encoding for fine-grained feature extraction, efficient spatial downsampling through learned attention and separate mechanisms for modeling frame-level and video-level relationships. Our model achieves comparable or superior performance to encoder-based approaches for open-ended video question answering on standard benchmarks. The fine-grained video question-answering evaluation demonstrates our model’s effectiveness, outperforming the encoder-based approaches Video-ChatGPT and Video-LLaVA in key aspects like correctness and temporal understanding. Extensive ablation studies validate our architectural choices and demonstrate the effectiveness of our spatio-temporal modeling approach while achieving 3-4$\times$ faster processing speeds than previous methods. Code is available at https://github.com/jh-yi/Video-Panda.
@@ 1 Introduction
Given these considerations, we ask: Is it possible to develop an efficient video-language model that bypasses the need for heavyweight encoders while maintaining state-of-the-art accuracy?. To address this question, we introduce a novel encoder-free video-language model that achieves remarkable performance with significantly reduced computational overhead. The key to our approach is a specialized spatio-temporal alignment block that effectively bridges visual and linguistic information without requiring image or video encoders.
@@ 2.2 Image-Language Models
More recently, an encoder-free approach has emerged, pioneered by Fuyu-8B [5], which processes image inputs directly through a decoder-only network. This architecture naturally handles high-resolution images with arbitrary aspect ratios, avoiding the constraints of traditional encoder-based systems. Recent advances in this direction include EVE [11], which introduces a novel training recipe focusing on unified decoder representation and enhanced visual recognition capabilities. These encoder-free models effectively address several key challenges: resolution constraints, deployment overhead, and capacity matching between vision and language components. Using minimal public training data, models like EVE [11] have demonstrated competitive performance with encoder-based approaches while requiring fewer computational resources.
@@ 2.3 Video-Language Models
Video-LLaMA [42] addressed multimodal integration by incorporating both visual and auditory content through a Video Q-former for temporal understanding and an Audio Q-former built on ImageBind for audio processing. Chat-UniVi [15] took a different approach by utilizing dynamic visual tokens to create a unified representation for both images and videos, employing a multi-scale representation to capture both high-level concepts and low-level details. Video-LLaVA [22] introduced a unified approach that aligns visual representations before projecting them into the language space. Combined with joint training on mixed image and video datasets, this method demonstrated strong performance across modalities. To the best of our knowledge, an encoder-free video-language model does not exist, representing a notable gap in the literature.
@@ 3.1 Spatio-Temporal Alignment Block (STAB)
$\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{P_{q}(\mathbf{Q})P_{k}(\mathbf{K})^{T}}{\sqrt{d}}\right)P_{v}(\mathbf{V})$ (4)
where $P_{q}$, $P_{k}$, and $P_{v}$ are learnable projection functions for query, key, and value transformations respectively. Then, for each window at position $(i,j)$ in frame $t$:
@@
yielding the downsampled representation $\mathbf{L}_{st}^{d}\in\mathbb{R}^{T\times h^{\prime}\times w^{\prime}\times d}$ where $h^{\prime}=h/2$ and $w^{\prime}=w/2$.
@@
$\mathbf{F}_{s,t}$ $=\text{Attn}(\mathbf{W}_{t},\mathbf{K}_{t},\mathbf{K}_{t})$ (10)
@@
$\mathbf{G}_{st}$ $=\text{Attn}(\mathbf{W}_{st},\mathbf{K}_{st},\mathbf{K}_{st})\in\mathbb{R}^{1\times d}$ (13)
@@
placing each frame’s conditioned representation $\mathbf{F}_{r,t}$ at the start of its corresponding token sequence. The final visual tokens are obtained by flattening $\mathbf{F}$ and projecting through an MLP to the language model’s embedding space:
@@ 4.1 Datasets and Implementation Details
Model #Param.(M) Inference time (ms)
VideoChatGPT [26] 307 171
Video-LLaVA [22] 425 125
Video-Panda 45 41
Table 3: Comparison of the number of parameters and inference time of the vision part.
@@ 6 Conclusion
This work introduces an efficient encoder-free approach to video-language understanding through our Spatio-Temporal Alignment Block (STAB). We reduce visual processing parameters by 6.5$\times$ or greater while maintaining competitive performance or outperforming comparable approaches in open-ended video question answering, obtaining 64.7% accuracy on MSVD-QA and 54.8% on MSRVTT-QA, as well as fine-grained video understanding. Additionally, STAB processes videos 3-4$\times$ faster than encoder-based methods. Our method shows promise for making video-language models more efficient, potentially benefiting longer videos and other tasks.
@@ Appendix A Detailed Hyperparameters
LR Schedule cosine decay
@@ Appendix B Additional Ablations
Table 10: Ablation study different teacher encoders.
@@ Appendix C Additional Dataset Details
    Correctness of Information: Evaluates the accuracy of generated text and its alignment with video content.
@@
Appendix D Additional Qualitative Ablations
@@
We present additional qualitative examples of our ablation studies in Figure 5, demonstrating Video-Panda’s effectiveness across various video understanding tasks. When using the complete training dataset in Stage 1 (left-top example), the model exhibits overfitting tendencies due to data imbalance, as evidenced by the example showing dog interactions—likely influenced by the disparity between dog (7,807) and cat (5,050) instances in the Valley dataset. The right-top example reveals that placing the LSD module before LSTE impairs cliff recognition due to early token downsampling and information loss, supporting our decision to maintain LSD placement after LSTE. Models using alternative approaches (average pooling, half resolution, or perceiver resampler) struggle with content recognition (e.g., cucumber, cat, pandas) compared to our learnable downsampling approach. Additionally, models using image-based teachers (CLIP and DINOv2) tend to make frame-specific predictions rather than considering global context, as demonstrated by their failure to recognize shredded potatoes across multiple frames.
Figure 5: Comparative analysis of Video-Panda qualitative ablation studies: The figure presents eight video examples with ground truth (GT) annotations and model predictions under different training configurations. The top row demonstrates: the effect of 702K training samples in stage 1, and the impact of performing Local Spatial Downsampling (LSD) before Local Spatial-Temporal Encoding (LSTE). The second row shows results from removing LSD while using: average pooling, half-resolution and perceiver resampler (third row left). The third row right and bottom row illustrate the effects of different teacher models: using Intern-Video and CLIP, and utilizing DINOv2 for knowledge distillation. Each example includes the original model prediction (yellow) and an ablated version (purple), highlighting how architectural and training choices affect Video-Panda’s ability to interpret dynamic visual scenes and answer questions accurately.
Appendix E EVE Baseline for Videos
@@
As the original EVE model [11] was designed for image processing, we conducted a fair comparison by re-training it (denoted as EVE* in Table 1) using identical video data (Valley-Pretrain-702K and Video-ChatGPT-100K datasets). For processing videos through EVE*, each frame was treated independently as a separate image, with CLIP-ViT-L/14 [32] serving as the teacher model for distillation. While this approach enables frame-level analysis, it neglects temporal relationships. In our implementation, we employ Learnable Selective Downsampling (LSD) to process video frames efficiently, reducing each frame to a consistent token count while preserving essential information. The resulting tokens are flattened into a single sequence, with special split tokens inserted between frame representations to maintain frame boundaries and enable temporal relationship learning.
Appendix F Broader Impact
@@
We introduce Video-Panda, an encoder-free Video Language Model for video understanding. Our model addresses key ethical and practical challenges in large-scale AI deployment. While many VLMs raise concerns about data bias, privacy, and computational costs, Video-Panda mitigates these issues through two key design choices: training exclusively on publicly available datasets and eliminating the need for a pretrained encoder. This approach not only reduces ethical concerns but also significantly lowers computational requirements and deployment costs, making the model more accessible and environmentally sustainable.

### v2

@@ Abstract
We present an efficient encoder-free approach for video-language understanding that achieves competitive performance while significantly reducing computational overhead. Current video-language models typically rely on heavyweight image encoders (300M-1.1B parameters) or video encoders (1B-1.4B parameters), creating a substantial computational burden when processing multi-frame videos. Our method introduces a novel Spatio-Temporal Alignment Block (STAB) that directly processes video inputs without requiring pre-trained encoders while using only 45M parameters for visual processing - at least a 6.5$\times$ reduction compared to traditional approaches. The STAB architecture combines Local Spatio-Temporal Encoding for fine-grained feature extraction, efficient spatial downsampling through learned attention and separate mechanisms for modeling frame-level and video-level relationships. Our model achieves comparable or superior performance to encoder-based approaches for open-ended video question answering on standard benchmarks. The fine-grained video question-answering evaluation demonstrates our model’s effectiveness, outperforming the encoder-based approaches Video-ChatGPT and Video-LLaVA in key aspects like correctness and temporal understanding. Extensive ablation studies validate our architectural choices and demonstrate the effectiveness of our spatio-temporal modeling approach while achieving 3-4$\times$ faster processing speeds than previous methods. Code is available at https://jh-yi.github.io/Video-Panda.
@@ 1 Introduction
Given these considerations, we ask: Is it possible to develop an efficient video-language model that bypasses the need for heavyweight encoders while maintaining state-of-the-art accuracy? To address this question, we introduce a novel encoder-free video-language model that achieves remarkable performance with significantly reduced computational overhead. The key to our approach is a specialized spatio-temporal alignment block that effectively bridges visual and linguistic information without requiring image or video encoders.
@@ 2.2 Image-Language Models
More recently, an encoder-free approach has emerged, pioneered by Fuyu-8B [5], which processes image inputs directly through a decoder-only network. This architecture naturally handles high-resolution images with arbitrary aspect ratios, avoiding the constraints of traditional encoder-based systems. Recent advances in this direction include EVE [11], which introduces a novel training recipe focusing on a unified decoder representation and enhanced visual recognition capabilities. These encoder-free models effectively address several key challenges: resolution constraints, deployment overhead, and capacity matching between vision and language components. Using minimal public training data, models like EVE [11] have demonstrated competitive performance with encoder-based approaches while requiring fewer computational resources.
@@ 2.3 Video-Language Models
Video-LLaMA [45] addressed multimodal integration by incorporating both visual and auditory content through a Video Q-former for temporal understanding and an Audio Q-former built on ImageBind for audio processing. Chat-UniVi [16] took a different approach by utilizing dynamic visual tokens to create a unified representation for both images and videos, employing a multi-scale representation to capture both high-level concepts and low-level details. ST-LLM [26], proposed as dynamic masking strategy which improved both the efficiency and robustness of the video VLM. Video-LLaVA [23] introduced a unified approach that aligns visual representations before projecting them into the language space. Combined with joint training on mixed image and video datasets, this method demonstrated strong performance across modalities. In this work, we address a notable gap in the literature and present the first encoder-free video-language model.
@@ 3.1 Spatio-Temporal Alignment Block (STAB)
$\text{Attn}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{P_{q}(\mathcal{Q})P_{k}(\mathcal{K})^{T}}{\sqrt{d}}\right)P_{v}(\mathcal{V})$ (4)
where $P_{q}$, $P_{k}$, and $P_{v}$ are learnable projection functions for query, key, and value transformations respectively. Then, for each window at position $(i,j)$ in frame $t$
@@
yields the downsampled representation $\mathbf{L}_{st}^{d}\in\mathbb{R}^{T\times h^{\prime}\times w^{\prime}\times d}$ where $h^{\prime}=h/2$ and $w^{\prime}=w/2$.
@@
$\mathbf{F}_{s,t}$ $=\text{Attn}(\mathbf{W}_{t},\mathbf{K}_{t},\mathbf{K}_{t})\,,$ (10)
@@
$\mathbf{G}_{st}$ $=\text{Attn}(\mathbf{W}_{st},\mathbf{K}_{st},\mathbf{K}_{st})\in\mathbb{R}^{1\times d}\,.$ (13)
@@
placing each frame’s conditioned representation $\mathbf{F}_{r,t}$ at the start of its corresponding token sequence. The final visual tokens are obtained by flattening $\mathbf{F}$ and projecting it through an MLP to the language model’s embedding space:
@@ 4.1 Datasets and Implementation Details
Model Vision Size (M) FLOPs (G) Latency (ms)
Video-ChatGPT [28] 307 8109.1 171
Video-LLaVA [23] 425 864.1 125
Video-Panda 45 105.5 41
Table 3: Comparison of the number of parameters, FLOPs, and latency of the vision part.
@@ 6 Conclusion
This work introduces an efficient encoder-free approach to video-language understanding through our Spatio-Temporal Alignment Block (STAB). We reduce the number of visual processing parameters by 6.5$\times$ or greater while maintaining competitive performance or outperforming comparable approaches in open-ended video question answering, obtaining 64.7% accuracy on MSVD-QA and 54.8% on MSRVTT-QA, as well as fine-grained video understanding. Additionally, STAB processes videos 3-4$\times$ faster than encoder-based methods. Our method shows promise for making video-language models more efficient, potentially benefiting longer videos and other tasks.
Acknowledgement
This work was supported by the Federal Ministry of Education and Research (BMBF) under grant no. 01IS22094A WEST-AI and the ERC Consolidator Grant FORHUE (101044724). Yanan Luo has been supported by the Chinese Scholarship Council (202108440041). For the computations involved in this research, we acknowledge EuroHPC Joint Undertaking for awarding us access to Leonardo at CINECA, Italy, through EuroHPC Regular Access Call - proposal No. EHPC-REG-2024R01-076.
@@ Appendix A Detailed Hyperparameters
LR Schedule cos. decay cos. decay cos. decay
@@ Appendix B Additional Ablations
Table 10: Ablation study on different teacher encoders.
@@ Appendix C Additional Dataset Details
    Correctness of Information: Evaluates accuracy of generated text and its alignment with video content.
@@
Appendix D Evaluation on Long Videos
To explore the potential of Video-Panda on long video benchmarks, we have evaluated our method on the EgoSchema [30] and Video-MME-M [13] datasets. The results presented in Table 11 confirm the results presented in the paper, i.e., we achieve similar or slightly better accuracy compared to Video-ChatGPT and Video-LLaVA, but require much less computational resources (Table 3).
Model Vision Size (M) EgoSchema VideoMME-M
Video-ChatGPT 307 34.2 36.0
Video-LLaVA 425 36.1 38.1
Video-Panda 45 36.4 37.9
Table 11: Results on the EgoSchema and VideoMME-M datasets.
Appendix E Additional Qualitative Ablations
@@
We present additional qualitative examples of our ablation studies in Figure 5, demonstrating Video-Panda’s effectiveness across various video understanding tasks. When using the complete training dataset in Stage 1 (left-top example), the model exhibits overfitting tendencies due to data imbalance, as evidenced by the example showing dog interactions—likely influenced by the disparity between dog (7,807) and cat (5,050) instances in the Valley dataset. The right-top example reveals that placing the LSD module before LSTE impairs cliff recognition due to early token downsampling and information loss. Models using alternative approaches (average pooling, half resolution, or perceiver resampler) struggle with content recognition (e.g., cucumber, cat, pandas) compared to our learnable downsampling approach. Additionally, models using image-based teachers (CLIP and DINOv2) tend to make frame-specific predictions rather than considering global context, as demonstrated by their failure to recognize shredded potatoes across multiple frames. We also provide additional qualitative examples on each dataset in Figure 6, Figure 7, and Figure 8.
Appendix F EVE Baseline for Videos
@@
As the original EVE model [11] was designed for image processing, we conducted a fair comparison by re-training it (denoted as EVE* in Table 1) using identical video data (Valley-702K and Video-ChatGPT-100K). For processing videos, each frame was treated independently as a separate image, with CLIP-ViT-L/14 [35] serving as the teacher model for distillation. While this approach enables frame-level analysis, it neglects temporal relationships. In our implementation, we employ Learnable Selective Downsampling (LSD) to process video frames efficiently, reducing each frame to a consistent token count while preserving essential information. The resulting tokens are flattened into a single sequence, with special split tokens inserted between frame representations to maintain frame boundaries and enable temporal relationship learning.
Appendix G Broader Impact
@@
We introduce Video-Panda, an encoder-free Video Language Model for video understanding. Our model addresses key practical challenges in large-scale AI deployment. While many VLMs raise concerns about data bias, privacy, and computational costs, Video-Panda mitigates these issues through two key design choices: training exclusively on publicly available datasets and eliminating the need for a pretrained encoder. This approach not only reduces ethical concerns but also significantly lowers computational requirements and deployment costs, making the model more accessible and environmentally sustainable.
Figure 5: Qualitative comparisons of different design choices of Video-Panda: The figure presents eight video examples with ground truth (GT) annotations and model predictions under different training configurations. The top row demonstrates the effect of 702K training samples in stage 1 (left) and the impact of performing Local Spatial Downsampling (LSD) before Local Spatial-Temporal Encoding (LSTE) (right). The second row shows results from removing LSD while using average pooling (left), half-resolution (right), and perceiver resampler (third row left). The third row right and fourth row illustrate the effects of different teacher models for knowledge distillation: CLIP (third row right), Intern-Video (left), and DINOv2 (right). Each example includes the original model prediction (yellow) and an ablated version (purple), highlighting how architectural and training choices affect Video-Panda’s ability to interpret dynamic visual scenes and answer questions. The qualitative examples are from the MSVD-QA dataset.
Figure 6: Qualitative examples from the MSRVTT-QA dataset.
Figure 7: Qualitative examples from the TGIF-QA dataset.
Figure 8: Qualitative examples from the ActivityNet-QA dataset.